OpenStudy (e.cociuba):

Solve for x. 4(x + 3)^2 + 5 = 10

OpenStudy (e.cociuba):

I need a bit help solving this problem. I'm afraid I forgot Algebra and it's been awhile since I have done it. I would appreciate some help. Thanks! :)

OpenStudy (goformit100):

For solving this do you know the conception of : Quadratic Equations ?

OpenStudy (e.cociuba):

Ok so this is where I got so far: 4(x + 3)^2 + 5 = 10 -5 -5 4(x + 3)^2= 5 This is where I forgot how to do it..

OpenStudy (e.cociuba):

K got it. Thanks :)

OpenStudy (anonymous):

\[4(x+3)^2+5=10\\\\ (x+3)^2= \frac{10}{4}\\ x+3 = \pm \sqrt {\frac{10}{ 4 }}\\ x+3 =\frac{ \pm \sqrt{5} }{ 2 }\\\\ x = -3 \pm\frac{ \sqrt{5} }{ 2 }\\\\ x = -3+\frac{\sqrt{5} }{ 2 }\ \ \ and \ \ \ \ x = -3-\frac{\sqrt{5} }{ 2 }\\ \]